Two forces, 3N and 4N act on a body in the direction due North and East respectively. Calculate the their equilibrium.
1
Expert's answer
2021-10-03T13:23:40-0400
"F=\sqrt{3^2+4^2}=5\ N\\\theta=\arctan{\frac{3}{4}}=37\degree\text{ N of E}"